OpenStudy (anonymous):

Which of the following ordered pairs lies on the graph of h(x) = -2x^2? (1, -4) (1, -2) (1, 4) Please help me!

OpenStudy (freckles):

did you plug in the points to see which will hold?

OpenStudy (anonymous):

I did. I believe it is (1, -2).

OpenStudy (freckles):

for example (1,-4) says we have h(1)=-2(1)^2=-4 is this equation true?

OpenStudy (freckles):

yes you are right -2(1)^2=-2 is totally true

OpenStudy (anonymous):

Okay, thank you. :)

OpenStudy (freckles):

np great job